# iKalibr-RGBD: Partially-Specialized Target-Free Visual-Inertial Spatiotemporal Calibration For RGBDs via Continuous-Time Velocity Estimation

### v1

@@ Abstract
Visual-inertial systems have been widely studied and applied in the last two decades, mainly due to their low cost and power consumption, small footprint, and high availability. Such a trend simultaneously leads to a large amount of visual-inertial calibration methods being presented, as accurate spatiotemporal parameters between sensors are a prerequisite for visual-inertial fusion. In our previous work, i.e., iKalibr, a continuous-time-based visual-inertial calibration method was proposed as a part of one-shot multi-sensor resilient spatiotemporal calibration. While requiring no artificial target brings considerable convenience, computationally expensive pose estimation is demanded in initialization and batch optimization, limiting its availability. Fortunately, this could be vastly improved for the RGBDs with additional depth information, by employing mapping-free ego-velocity estimation instead of mapping-based pose estimation. In this paper, we present the continuous-time ego-velocity estimation-based RGBD-inertial spatiotemporal calibration, termed as iKalibr-RGBD, which is also targetless but computationally efficient. The general pipeline of iKalibr-RGBD is inherited from iKalibr, composed of a rigorous initialization procedure and several continuous-time batch optimizations. The implementation of iKalibr-RGBD is open-sourced at (https://github.com/Unsigned-Long/iKalibr) to benefit the research community.
@@ I Introduction
Visual-inertial systems, generally composed of vision sensors and inertial measurement units (IMUs), have been applied in a wide range of robotic tasks, mainly for ego-motion estimation [1, 2] and navigation [3, 4]. With constantly emerging new vision sensors, such as depth-aware RGBD and change-aware event cameras, visual-inertial applications become more extensive and ubiquitous. For such visual-inertial sensor suites, spatiotemporal calibration is required to determine unique spatial and temporal frameworks for accurate and consistent data fusion.
Visual-inertial spatiotemporal calibration, regardless of the type of vision sensor integrated in the suite, could usually be achieved by employing existing well-studied calibration methods [5, 6, 7, 8] orienting to the conventional optical cameras (capturing color or gray images), since conventional optical images can be generally provided by these mechanism-disparate vision sensors (mainly benefiting from supports of manufacturers). Such a calibration route is also followed by our previous work iKalibr [9], whose visual-inertial spatiotemporal calibration exactly focuses on optical vision sensors. As a targetless, continuous-time, and open-source calibrator, iKalibr brings considerable convenience to spatiotemporal calibration. However, significant computation cost exists in iKalibr, especially in visual-inertial calibration, as the computationally expensive mapping-based structure from motion (SfM) [10] is required for each optical camera in calibration. Fortunately, this can be avoided for RGBD cameras with available depth information, through partially specialized visual modeling, which is exactly the RGBD-inertial spatiotemporal calibration this work presents.
@@ II-A Visual-Inertial Calibration
Artificial visual targets, such as chessboards and April Tags, are commonly utilized in the early target-based visual-inertial calibration works for efficient and accurate data association. Based on the Kalman filter, Mirzaei et al. [13] presented a checker-board-based visual-inertial extrinsic calibration method. Similarly relying on chessboards but employing continuous-time-based estimation, Furgale et al. [5] proposed the well-known Kalibr for both spatial (extrinsic) and temporal calibration of a visual-inertial sensor suite, which is the first continuous-time-based spatiotemporal calibration. This work was subsequently extended by Huai et al. [14] to support rolling shutter cameras besides global shutter ones. In addition to the above target-based calibration, target-free methods are also studied, most of which are online ones integrated into real-time visual-inertial estimators. Focusing on online temporal determination, Li et al. [15] presented an extended Kalman filter (EKF)-based visual-inertial odometry, where the time offset to be calibrated is treated as a part of the state vector. Besides temporal calibration, online extrinsic determination is also studied in [16] by Hartzer et al., which is also EKF-based but supports multiple cameras. More advancedly, Huai et al. [17] presented an EKF-based visual-inertial odometry with online extrinsic, intrinsic, and temporal calibration. In our previous work [9], we proposed a continuous-time-based and targetless spatiotemporal calibration method for integrated inertial systems, one of whose modules is visual-inertial calibration. Note that the above-discussed calibration methods orient to normal optical cameras, but can also apply to RGBD cameras, as RGB or gray images can be generally provided by RGBDs.
@@ III Preliminaries
This section details the notations, sensor models, as well as the B-spline-based continuous-time time-varying state representation used in this paper.
@@ V Real-World Experiment
Sufficient real-world experiments were carried out to verify the feasibility of the proposed iKalibr-RGBD and evaluate its performance in terms of calibration repeatability, accuracy, consistency, and computation consumption.
@@ V-B Spatiotemporal Convergence
The spatiotemporal convergence performance of the proposed iKalibr-RGBD is shown in Fig. 6, where spatiotemporal parameters of two cameras are with respect to the IMU. All spatiotemporal parameters at the beginning are set to zeros or identities. For better readers’ understanding, the horizontal axis is labeled by indexes of corresponding sections. As can be seen, in the initialization procedure (Section IV-B, IV-C, and IV-D), the extrinsic rotation, time offset, and extrinsic translation are initialized sequentially, where two B-splines are also recovered. Subsequently, two batch optimizations are conducted (Section IV-E), which refines all states to the global optimal ones. These results demonstrate that iKalibr-RGBD could effectively and efficiently estimate spatiotemporal parameters and reach the final convergence.
@@ V-C Accuracy and Repeatability Comparison
Param. iKalibr-RGBD iKalibr [9] Kalibr* [5] CAD*
Cam-Left $\hat{p}_{x}$ 11.76$\pm$0.54 12.20$\pm$0.42 14.16 11.32
$\hat{p}_{y}$ 8.09$\pm$0.35 8.35$\pm$0.11 7.36 8.25
$\hat{p}_{z}$2.03$\pm$0.51 -1.50$\pm$0.14 -1.47 -1.45
$\hat{\theta}_{r}$ 90.46$\pm$0.11 90.55$\pm$0.03 90.97 90.00
$\hat{\theta}_{p}$179.53$\pm$0.04 -179.51$\pm$0.01 -179.54 -180.00
$\hat{\theta}_{y}$ 90.10$\pm$0.15 90.07$\pm$0.02 90.46 90.00
$\hat{\tau}$ 9.59$\pm$0.75 9.66$\pm$0.70 8.46
Cam-Right $\hat{p}_{x}$ 12.02$\pm$0.47 12.30$\pm$0.46 13.97 11.32
$\hat{p}_{y}$8.53$\pm$0.39 -8.72$\pm$0.11 -9.77 -8.75
$\hat{p}_{z}$1.80$\pm$0.58 -1.42$\pm$0.21 -1.37 -1.45
$\hat{\theta}_{r}$ 90.22$\pm$0.08 90.25$\pm$0.05 90.56 90.00
$\hat{\theta}_{p}$ 179.02$\pm$0.04 179.05$\pm$0.02 179.16 180.00
$\hat{\theta}_{y}$ 90.42$\pm$0.12 90.27$\pm$0.02 90.36 90.00
$\hat{\tau}$ 9.58$\pm$0.78 9.64$\pm$0.70 8.39
    Extrinsic translations in $(cm)$, extrinsic Euler angles in $(deg)$, and time offsets in $(ms)$. All parameters are with respect to the IMU.
@@
The calibration accuracy and repeatability of iKalibr-RGBD are subsequently evaluated. Table I summarizes the calibration results of iKalibr-RGBD in the five scenarios mentioned above, where spatiotemporal parameters from Kalibr [5] and computer aided design (CAD) provided by VECtor Benchmark [28] are also listed. To evaluate the repeatability, the visual-inertial calibration module of iKalibr [9] orienting to optical cameras is also considered in the experiments for a reliable comparison. It could be found that the proposed iKalibr-RGBD achieves comparable calibration results with its predecessor iKalibr. The STDs of extrinsic rotation, extrinsic translation, and time offset are within $0.15\;deg$, $0.6\;cm$, and $0.8\;ms$, respectively. Although iKalibr-RGBD performs slightly worse than iKalibr in terms of repeatability (see the STDs), mainly on extrinsic translation, due to employing non-primitive visual observations, it has obvious advantages in terms of computation efficiency, see Section V-E and Table II.
@@ V-D Consistency Evaluation
The calibration consistency is subsequently evaluated based on the by-products from spatiotemporal determination. Based on the estimated B-splines, we project the world-frame gravity vector and linear velocities of tracked landmarks with respect to the camera onto the image plane, see Fig. 7. Estimates-derived and optical-flow-derived pixel velocities of features, i.e., $\hat{\dot{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}$ from (2) and $\tilde{\dot{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}$ from (17), are also plotted to qualitatively evaluate consistency. As can be seen, the direction of estimated gravity is consistent well with the structures of surroundings. Meanwhile, estimates-derived pixel velocities are well matched with the ones derived by optical flow. To quantitatively evaluate their difference (i.e., visual residual), the two-dimensional visual optical flow residuals $r_{c}\left(\dot{{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}\right)$ after final spatiotemporal convergence are statistically analyzed and depicted in Fig. 8. It could be found that each dimension of errors obeys zero-mean Gaussian distribution, and STDs are generally within $20\;pixels/sed$. These results show the calibration consistency iKalibr-RGBD is capable of.
@@ V-E Computation Consumption Comparison
    Time consumption statistics are based on spatiotemporal calibration of a single-camera (left camera) single-IMU sensor configuration. Data used in calibration lasts $30\;sed$ (about 900 images).
@@
The most impressive point of the proposed iKalibr-RGBD is its efficient computation and low memory usage, compared with the visual-inertial calibration of its predecessor iKalibr, see Table II. Such a high performance mainly comes from the mapping-free ego-velocity estimation iKalibr-RGBD employed, which does not require time-consuming SfM [10] for visual-inertial initialization and subsequent mapping-based continuous-time trajectory optimization iKalibr adopted.
@@ Acknowledgments
The calibration data acquisition is performed on the GREAT software, developed by the GREAT Group, School of Geodesy and Geomatics (SGG), Wuhan University.

### v2

@@ Abstract
Visual-inertial systems have been widely studied and applied in the last two decades (from the early 2000s to the present), mainly due to their low cost and power consumption, small footprint, and high availability. Such a trend simultaneously leads to a large amount of visual-inertial calibration methods being presented, as accurate spatiotemporal parameters between sensors are essential for visual-inertial fusion. In our previous work, i.e., iKalibr, a continuous-time-based visual-inertial calibration method was proposed as a part of one-shot multi-sensor resilient spatiotemporal calibration. While requiring no artificial calibration target brings considerable convenience, computationally expensive pose estimation is demanded in initialization and batch optimization, limiting its availability. Fortunately, this can be vastly improved for the RGBDs with additional depth information, by employing mapping-free ego-velocity estimation instead of mapping-based pose estimation. In this paper, we present an ego-velocity-estimation-based RGBD-inertial spatiotemporal calibrator, termed as iKalibr-RGBD, which is also targetless and continuous-time-based, but computationally efficient. The general pipeline of iKalibr-RGBD is inherited from iKalibr, composed of a rigorous initialization procedure and several continuous-time batch optimizations. The implementation of iKalibr-RGBD is open-sourced at (https://github.com/Unsigned-Long/iKalibr) to benefit the research community.
@@ I Introduction
Visual-inertial systems, generally composed of vision sensors and inertial measurement units (IMUs), have been applied in a wide range of robotic tasks, mainly for ego-motion estimation [1, 2] and navigation [3, 4]. With constantly emerging new vision sensors, such as depth-aware RGBD and change-aware event cameras, visual-inertial applications become more extensive and ubiquitous. For such visual-inertial sensor suites, accurate spatiotemporal calibration is crucial for reliable data fusion.
Visual-inertial spatiotemporal calibration, regardless of the type of vision sensor integrated in the suite, could usually be achieved by employing existing well-studied calibration methods [5, 6, 7, 8] orienting to the conventional optical cameras (capturing color or gray images), since conventional optical images can be generally provided by these mechanism-disparate vision sensors (mainly benefiting from supports of manufacturers). Such a calibration route is also followed by our previous work iKalibr [9], whose visual-inertial spatiotemporal calibration exactly focuses on optical vision sensors. As a targetless, continuous-time, and open-source calibrator, iKalibr provides significant benefits to spatiotemporal calibration. However, significant computation cost exists in iKalibr, especially in visual-inertial calibration, as the computationally expensive mapping-based structure from motion (SfM) [10] is required for each optical camera in calibration. Fortunately, this can be avoided for RGBD cameras with available depth information, through partially specialized visual modeling, which is exactly the RGBD-inertial spatiotemporal calibration this work presents.
@@ II-A Visual-Inertial Calibration
Artificial visual targets, such as chessboards and April Tags, are commonly utilized in the early target-based visual-inertial calibration works for efficient and accurate data association. Based on the Kalman filter, Mirzaei et al. [13] presented a checker-board-based visual-inertial extrinsic calibration method. Similarly relying on chessboards but employing continuous-time-based estimation, Furgale et al. [5] proposed the well-known Kalibr for both spatial (extrinsic) and temporal calibration of a visual-inertial sensor suite, which is the first continuous-time-based spatiotemporal calibration. This work was subsequently extended by Huai et al. [14] to support rolling shutter cameras besides global shutter ones. In addition to the above target-based calibration, target-free methods are also studied, most of which are online ones integrated into real-time visual-inertial estimators. Focusing on online temporal determination, Li et al. [15] presented an extended Kalman filter (EKF)-based visual-inertial odometry, where the time offset to be calibrated is treated as a part of the state vector. Besides temporal calibration, online extrinsic determination is also studied in [16] by Hartzer et al., which is also EKF-based but supports multiple cameras. More recently, Huai et al. [17] presented an EKF-based visual-inertial odometry with online extrinsic, intrinsic, and temporal calibration. In our previous work [9], we proposed a continuous-time-based and targetless spatiotemporal calibration method for integrated inertial systems, one of whose modules is visual-inertial calibration. Note that the above-discussed calibration methods orient to normal optical cameras, but can also apply to RGBD cameras, as RGB or gray images can be generally provided by RGBDs.
@@ III Preliminaries
This section details the notations, sensor models, and the B-spline-based continuous-time time-varying state representation used in this paper.
@@ V Real-World Experiment
A sufficient number of real-world experiments were carried out to verify the feasibility of the proposed iKalibr-RGBD and evaluate its performance in terms of calibration repeatability, accuracy, consistency, and computation consumption.
@@ V-B Spatiotemporal Convergence
The spatiotemporal convergence performance of the proposed iKalibr-RGBD is shown in Fig. 6, where spatiotemporal parameters of two cameras are with respect to the IMU. All spatiotemporal parameters at the beginning are set to zeros or identities. For a better understanding by readers, the horizontal axis is labeled by indexes of corresponding sections. As can be seen, in the initialization procedure (Section IV-B, IV-C, and IV-D), the extrinsic rotation, time offset, and extrinsic translation are initialized sequentially, where two B-splines are also recovered. Subsequently, two batch optimizations are conducted (Section IV-E), which refines all states to the global optimal ones. These results demonstrate that iKalibr-RGBD could effectively and efficiently estimate spatiotemporal parameters and reach the final convergence.
@@ V-C Accuracy and Repeatability Comparison
Param. iKalibr-RGBD iKalibr [9] Kalibr [5] CAD
Cam-L To IMU $\hat{p}_{x}$ 12.27$\pm$0.47 12.20$\pm$0.42 14.16 11.32
$\hat{p}_{y}$ 8.32$\pm$0.22 8.35$\pm$0.11 7.36 8.25
$\hat{p}_{z}$1.79$\pm$0.23 -1.50$\pm$0.14 -1.47 -1.45
$\hat{\theta}_{r}$ 90.54$\pm$0.05 90.55$\pm$0.03 90.97 90.00
$\hat{\theta}_{p}$179.51$\pm$0.02 -179.51$\pm$0.01 -179.54 -180.00
$\hat{\theta}_{y}$ 90.83$\pm$0.08 90.07$\pm$0.02 90.46 90.00
$\hat{\tau}$ 9.62$\pm$0.80 9.66$\pm$0.70 8.46
Cam-R To IMU $\hat{p}_{x}$ 12.34$\pm$0.41 12.30$\pm$0.46 13.97 11.32
$\hat{p}_{y}$8.48$\pm$0.24 -8.72$\pm$0.11 -9.77 -8.75
$\hat{p}_{z}$1.71$\pm$0.30 -1.42$\pm$0.21 -1.37 -1.45
$\hat{\theta}_{r}$ 90.30$\pm$0.02 90.25$\pm$0.05 90.56 90.00
$\hat{\theta}_{p}$ 179.05$\pm$0.03 179.05$\pm$0.02 179.16 -180.00
$\hat{\theta}_{y}$ 90.59$\pm$0.09 90.27$\pm$0.02 90.36 90.00
$\hat{\tau}$ 9.61$\pm$0.72 9.64$\pm$0.70 8.39
Cam-R To Cam-L $\hat{p}_{x}$ 16.80$\pm$0.26 17.07$\pm$0.06 17.14 17.00
$\hat{p}_{y}$ 0.06$\pm$0.12 0.05$\pm$0.09 0.03 0.00
$\hat{p}_{z}$0.19$\pm$0.29 -0.17$\pm$0.07 -0.48 0.00
$\hat{\theta}_{r}$0.22$\pm$0.05 -0.30$\pm$0.04 -0.09 0.00
$\hat{\theta}_{p}$ 0.35$\pm$0.08 0.20$\pm$0.03 0.42 0.00
$\hat{\theta}_{y}$ 1.43$\pm$0.02 1.43$\pm$0.02 1.29 0.00
$\hat{\tau}$0.01$\pm$0.04 -0.01$\pm$0.02 -0.07
    Extrinsic translations in $(cm)$, extrinsic Euler angles in $(deg)$, and time offsets in $(ms)$.
@@
The calibration accuracy and repeatability of iKalibr-RGBD are subsequently evaluated. Table I summarizes the calibration results of iKalibr-RGBD in the five scenarios mentioned above, where spatiotemporal parameters from Kalibr [5] and computer aided design (CAD) provided by VECtor Benchmark [28] are also listed. To evaluate the repeatability, the visual-inertial calibration module of iKalibr [9] orienting to optical cameras is also considered in the experiments for a reliable comparison. It could be found that the proposed iKalibr-RGBD achieves comparable calibration results with its predecessor iKalibr. The STDs of extrinsic rotation, extrinsic translation, and time offset are within $0.1\;deg$, $0.5\;cm$, and $0.8\;ms$, respectively. Although iKalibr-RGBD performs slightly worse than iKalibr and Kalibr in terms of repeatability (see the STDs) and accuracy, mainly on scale-related extrinsic translation, due to employing non-primitive visual observations and the lack of constraints from explicit priori geometric scale compared with Kalibr, it has obvious advantages in terms of computation efficiency, see Section V-E and Table II.
@@ V-D Consistency Evaluation
The calibration consistency is subsequently evaluated based on the by-products from spatiotemporal determination. Based on the estimated B-splines, we project the world-frame gravity vector and linear velocities of tracked landmarks with respect to the camera onto the image plane, see Fig. 7. Estimates-derived and optical-flow-derived pixel velocities of features, i.e., $\hat{\dot{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}$ from (2) and $\tilde{\dot{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}$ from (17), are also plotted to qualitatively evaluate consistency. As can be seen, the direction of estimated gravity is consistent well with the structures of surroundings. Meanwhile, estimates-derived pixel velocities are well matched with the ones derived by optical flow. To quantitatively evaluate their difference (i.e., visual residual), the two-dimensional visual optical flow residuals $r_{c}\left(\dot{{\boldsymbol{\mathrm{f}}}}_{l_{m}}^{c_{n}}\right)$ after final spatiotemporal convergence are statistically analyzed and depicted in Fig. 8. It could be found that each dimension of errors obeys zero-mean Gaussian distribution, and STDs are generally within $20\;pixels/sec$. These results show the calibration consistency iKalibr-RGBD is capable of.
@@ V-E Computation Consumption Comparison
Kalibr Corner Extraction+Opt. 0.252+1.272=1.524
    Time consumption statistics are based on spatiotemporal calibration of a single-camera (left camera) single-IMU sensor configuration. Data used in calibration lasts $30\;sec$ (about 900 images).
@@
The most impressive point of the proposed iKalibr-RGBD is its efficient computation, compared with the visual-inertial calibration of its predecessor iKalibr and Kalibr, see Table II. It can be found that iKalibr-RGBD outperforms iKalibr and Kalibr on computational efficiency. Such a high performance mainly comes from the mapping-free ego-velocity estimation iKalibr-RGBD employed, which does not require time-consuming SfM [10] for visual-inertial initialization and subsequent mapping-based continuous-time trajectory optimization iKalibr adopted. Meanwhile, as the image acquisition frequency increases, such a difference in computational consumption will be more significant.
@@ Acknowledgments
The optimization for code implementation of iKalibr-RGBD was assisted partially by the Supercomputing Center and the GREAT Group of Wuhan University.